OpenStudy (anonymous):

Thirteen plums weigh as much as two apples and one pear. Four plums and one apple have the same weight as one pear. How many plums have the weight of one pear?

OpenStudy (anonymous):

we can put this in an equation thing firstly. I will say that refers to plums a to apple and k to pears

OpenStudy (anonymous):

13p=2a+1k 1k=4p+1a

OpenStudy (anonymous):

thats what i did but i dont know what to do from there

OpenStudy (anonymous):

hmm, try solving it

OpenStudy (anonymous):

* let me try solving it that's what I meant

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok, got it

OpenStudy (anonymous):

ok, so we know that 13p=2a+1k and that 4p+1a=1k first thing I did was to get both equation to be solving for the same thing. so I moved 2a on the first equation to the left and have 13p-2a=1k and since the other equation is 4p+1a=1k I know that 13p-2a=4p+1a. See how I got that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you get an idea what to do next?

OpenStudy (anonymous):

yes thank you so much!

OpenStudy (anonymous):

glad I did :)